OpenStudy (anonymous):

[x(x+1)]/2 = (2x-3)/3 +1 I will give a medal to someone who can explain the steps, thanks!

OpenStudy (anonymous):

\[\frac{ x(x+1) }{ 2 }=\frac{ 2x-3 }{ 3 }+1\] I will give a medal to someone who can explain the steps, thanks!

OpenStudy (anonymous):

solve for \(x\) ?

OpenStudy (anonymous):

probably easiest thing to do is get rid of annoying denominators by multiplying both sides of the equation by \(6\) to start with \[3x(x-1)=2(2x-3)+6\] now at least you have no denominators

OpenStudy (anonymous):

you good from there? multiply out, and solve the resulting quadratic equation

OpenStudy (anonymous):

I think so, thanks!

OpenStudy (anonymous):

k let me know if you need other help, but it is now much easier without denominators yw